OpenStudy (anonymous):

How does this −1/5∑n=0∞(x∗xn))/5n(n+1) turn into −∑n=1∞(xn)/(n5n) I got to the first part, and looked into a solution manual and it showed the second part.. I am confused on how it came about.

OpenStudy (anonymous):

ew it copied wrong \[-(1/5)\sum_{n=0}^{\infty} (x*x^n)/(5^n(n+1))\]\[= -\sum_{n=1}^{\infty} (x^n)/(n5^n)\]